# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Results & discussion - The entanglement between the optical and mechanical subsystems is quantified by the smallest symplectic eigenvalue, $\nu_{-}$, of the partially transposed covariance matrix [40, 41]. With our normalization, $\nu_{-}=1$ if the two oscillators are in uncorrelated vacuum states, and the presence of entanglement is signaled by $\nu_{-}<1$. Fig. 3 reports the values of $\nu_{-}$ calculated from the covariance matrix $\mathbf{V}$ directly reconstructed through $\mathbf{A}$, as a function of the width $\Gamma_{\scriptscriptstyle{\xi}}$ of $\xi(\omega)$. Entanglement emerges when the bandwidth $\Gamma_{\scriptscriptstyle{\xi}}$ of the optical mode becomes sufficiently large to encompass the full spectral response of the mechanical bright mode [17], characterized by two peaks at $\sim\Omega_{\scriptstyle{x}}$ and $\sim\Omega_{\scriptstyle{y}}$ broadened by optomechanical interaction. The minimum value of $\nu_{-}$, corresponding to maximal entanglement, is achieved around $\Gamma_{\scriptscriptstyle{\xi}}/2\pi=40\,$kHz and reaches $\nu_{-}=0.931\pm 0.015$. Here, the quoted statistical uncertainty represents one standard deviation obtained from multiple measurements, while an additional systematic uncertainty of $\pm 0.02$ arises from independent estimates of $\eta$ and $\kappa$.
@@
Entanglement is commonly quantified by the logarithmic negativity $E_{N}$, defined as $E_{N}=max\{0,-\log_{2}(\nu_{-})\}$ [42, 43]. At its maximum, we obtain $E_{N}=0.10\pm 0.03$, including both the statistical and the systematic uncertainty.
@@
Figure 3: Optimization of the optical mode width. The smallest symplectic eigenvalue $\nu_{-}$ of the partially trasposed covariance matrix $\mathbf{V}$ is plotted as a function of mode width $\Gamma_{\xi}$. The data points with error bars (indicating one SD over about 10 independent data acquisitions) are obtained from $\mathbf{V}$ directly reconstructed from the experimental correlation matrix. The green shaded area derives from $\mathbf{V}$ computed with our optomechanical model, using the parameters that best fit the heterodyne spectra. The width of this band reflects systematic uncertainties in the model parameters (mainly the efficiency $\eta$), with negligible contribution from statistical uncertainty. The gray shaded region highlights the threshold above which the mechanical and optical subsystems become separable.
@@
Fig. 3 also reports the values of $\nu_{-}$ obtained by estimating $\mathbf{V}$ from the full optomechanical model using all the fitted parameters, including the heating rates. The model reproduces the overall trend of the values extracted directly from the experimental covariance matrix, but consistently yields lower values of $\nu_{-}$. We attribute this discrepancy to slow fluctuations of the system parameters during the acquisition of the full time series, most notably of the relative phases $\phi_{A}$ and $\phi_{B}$, which are not fully captured by the stationary description. To mitigate these effects, the data analysis is restricted to time intervals over which the fitted parameters remain approximately constant. In practice, each independent estimate of the correlation matrix $\mathbf{A}$ — and hence of the covariance matrix $\mathbf{V}$ and the symplectic eigenvalue $\nu_{-}$ — is obtained by averaging over typically 2000 consecutive time windows of $10$ ms. Residual parameter fluctuations during these 20 seconds segments nevertheless lead to an effective mixing of slightly different Gaussian states, which results in a reduction of the observable entanglement. As a consequence, the entanglement actually present in the system is more faithfully captured by estimating $\mathbf{V}$ through the optomechanical model evaluated with the fitted parameters, whereas the values of $\nu_{-}$ directly inferred from the data should be regarded as a conservative estimate of the entanglement that can be effectively harnessed under the present level of system stability.
@@
We also notice that we have defined the optical mode using a rectangular spectral filter. Although convenient, this choice is not optimal. A suitably tailored spectral mode could maximize the fidelity between the optical and mechanical subsystems and thereby enhance the observable entanglement.
@@
The results discussed so far were obtained for detunings satisfying $\Delta_{A}\simeq\Delta_{B}\simeq\Omega_{\scriptstyle{b}}$. This condition, however, is not essential, and a finite degree of entanglement is present even when it is relaxed. Indeed, we repeated the experiment while varying the detuning, keeping only the condition $\Delta_{A}\simeq\Delta_{B}$. Fig. 4 shows the symplectic eigenvalue $\nu_{-}$ derived from the covariance matrix $\mathbf{V}$ reconstructed directly from the measured correlation matrix $\mathbf{A}$, together with the corresponding values obtained by computing $\mathbf{V}$ with the optomechanical model using parameters fitted to the same experimental data. Entanglement persists over a detuning range exceeding 40 kHz, consistent with the expectation that it is maintained within a bandwidth on the order of the cavity linewidth $\kappa$. This robustness implies that stable opto-mechanical entanglement can be achieved without fine control of the detuning.
@@
Figure 4: Symplectic eigenvalue versus cavity detuning. Experimental data points shown by blue markers with error bars (SEM over about 10 independent data acquisitions) are obtained from the covariance matrix directly reconstructed from the measured correlations. Green markers are extracted instead from the same experimental data through the optomechanical model, using the parameters that best fit the heterodyne spectra. In this case the uncertainty is dominated by the systematic error on the detection efficiency $\eta$. In these measurements, we set $\Delta_{A}\simeq\Delta_{B}=|\Delta|$. Intracavity entanglement is inferred from the symplectic eigenvalues associated to the covariance matrix calculated, within the same model and parameter set, for the joint state of the bright mode and the intracavity field (red points, with error bars again set by the systematic uncertainty on $\eta$). The green and red shaded regions are obtained by taking the extremal model predictions generated by the full set of parameters extracted from all data sets (at all detunings). These bands therefore illustrate the overall detuning dependence of the entanglement in our system.
@@
An important and nontrivial question is whether entanglement is also present inside the cavity, between the intracavity mode of field $B$ and the mechanical oscillator. This is reasonable but not obvious, since the intracavity field mode is uniquely determined by the cavity boundary conditions and cannot be tailored as in the case of the propagating field. Although intracavity observables cannot be accessed directly, the same optomechanical model that quantitatively reproduces the measured output correlations allows us to reconstruct the corresponding intracavity covariance matrix from the fitted parameters. The resulting symplectic eigenvalue $\nu_{-}$, shown in Fig. 4 as a function of the detuning, confirms the presence of intracavity entanglement, albeit at a reduced level compared to that of the propagating mode. The minimum inferred value of $\nu_{-}$ for the intracavity variables is $\nu_{-}=0.885\pm 0.003$, corresponding to a logarithmic negativity of $E_{N}=0.176\pm 0.005$, while for the propagating optical mode the same dataset gives $\nu_{-}=0.860\pm 0.013$ ($E_{N}=0.218\pm 0.023$).
@@
In conclusion, we have demonstrated the generation of quantum entanglement between the center-of-mass motion of a levitated nanosphere and an optical field, extending beyond the interaction region and persisting in the propagating wave. By combining coherent scattering in an optical cavity with high efficiency heterodyne detection, we reconstruct the full set of optical–mechanical correlations and observe the violation of separability bounds under experimentally accessible conditions. These results establish levitated optomechanical systems as a versatile platform for achieving and exploiting light–matter entanglement. The ability to produce it at room temperature, without requiring ultra-cryogenic environments or fine-tuned parameters, represents a significant step toward practical macroscopic quantum technologies. A distinctive advantage of levitated systems is the availability of additional, weakly coupled mechanical degrees of freedom beyond center-of-mass motion, including rotational and librational modes [44, 45, 46, 47]. Access to these additional mechanical degrees of freedom, combined with the prospect of trapping and simultaneously cooling multiple levitated nanospheres within a single optical cavity [48, 49], could enable the generation of multipartite and multimode entangled states mediated by a common optical field [50], while opening new avenues for scalable light–matter interfaces with enhanced sensitivity and control. More broadly, the combination of strong isolation, tunable interactions, and the ability to distribute quantum correlations via propagating fields positions levitated optomechanics as a promising setting for future investigations of fundamental physics. In particular, these systems suggest a realistic path toward experimental tests of whether gravity can mediate quantum correlations, by enabling the preparation and control of nonclassical states of increasingly massive objects [51, 52]. Such experiments could provide new insights into the quantum nature of gravity and into possible modifications of quantum mechanics at macroscopic scales.
@@ Acknowledgments
The authors would like to acknowledge useful discussions with D. Vitali and P. F. Barker. We acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI and by the European Commission-EU under the Infrastructure I-PHOQS “Integrated Infrastructure Initiative in Photonic and Quantum Sciences ” [IR0000016, ID D2B8D520, CUP D2B8D520].

### v2

@@ Abstract
Results & discussion - The entanglement between the optical and mechanical subsystems is quantified by the smallest symplectic eigenvalue, $\nu_{-}$, of the partially transposed covariance matrix [19, 45]. With our normalization, $\nu_{-}=1$ if the two oscillators are in uncorrelated vacuum states, and the presence of entanglement is signaled by $\nu_{-}<1$. Fig. 3 reports the values of $\nu_{-}$ calculated from the covariance matrix $\mathbf{V}$ directly reconstructed through $\mathbf{A}$, as a function of the width $\Gamma_{\scriptscriptstyle{\xi}}$ of $\xi(\omega)$. Entanglement emerges when the bandwidth $\Gamma_{\scriptscriptstyle{\xi}}$ of the optical mode becomes sufficiently large to encompass the full spectral response of the mechanical bright mode [21], characterized by two peaks at $\sim\Omega_{\scriptstyle{x}}$ and $\sim\Omega_{\scriptstyle{y}}$ broadened by optomechanical interaction. The minimum value of $\nu_{-}$, corresponding to maximal entanglement, is achieved around $\Gamma_{\scriptscriptstyle{\xi}}/2\pi=35\,$kHz and reaches $\nu_{-}=0.918\pm 0.029$. Here, the quoted statistical uncertainty represents one standard deviation obtained from multiple measurements, while an additional systematic uncertainty of $\pm 0.02$ arises from the independent estimates of $\eta$ and $\kappa$.
@@
Entanglement is commonly quantified by the logarithmic negativity $E_{N}$, defined as $E_{N}=max\{0,-\log_{2}(\nu_{-})\}$ [49, 36]. At its maximum, we obtain $E_{N}=0.12\pm 0.04$, including both the statistical and the systematic uncertainty.
@@
Figure 3: Optimization of the optical mode width. The smallest symplectic eigenvalue $\nu_{-}$ of the partially trasposed covariance matrix $\mathbf{V}$ is plotted as a function of the mode width $\Gamma_{\xi}$. The data points with error bars (indicating the SEM over about 10 independent data acquisitions) are obtained from $\mathbf{V}$ directly reconstructed from the experimental correlation matrix. The green shaded area derives from $\mathbf{V}$ computed with our optomechanical model, using the parameters that best fit the heterodyne spectra. The width of this band reflects systematic uncertainties in the model parameters (mainly the efficiency $\eta$), with negligible contribution from statistical uncertainty. The gray shaded region highlights the threshold above which the mechanical and optical subsystems become separable.
@@
Fig. 3 also reports the values of $\nu_{-}$ obtained by estimating $\mathbf{V}$ from the full optomechanical model using all the fitted parameters, including the heating rates. The model reproduces the overall trend of the values extracted directly from the experimental covariance matrix, but consistently yields lower values of $\nu_{-}$, with a minimum of $\nu_{-}=0.884\pm 0.011$ ($E_{N}=0.177\pm 0.019$). We attribute this discrepancy to slow fluctuations of the system parameters during the acquisition of the full time series, most notably of the relative phases $\phi_{A}$ and $\phi_{B}$, which are not fully captured by the stationary description. To mitigate these effects, the data analysis is restricted to time intervals over which the fitted parameters remain approximately constant. In practice, each independent estimate of the correlation matrix $\mathbf{A}$ — and hence of the covariance matrix $\mathbf{V}$ and the symplectic eigenvalue $\nu_{-}$ — is obtained by averaging over typically 2000 consecutive time windows of $10$ ms. Residual parameter fluctuations during this 20 s acquisition time nevertheless mix slightly different Gaussian states and reduce the observable entanglement. As a consequence, the entanglement actually present in the system is more faithfully captured by estimating $\mathbf{V}$ through the optomechanical model evaluated with the fitted parameters, whereas the values of $\nu_{-}$ directly inferred from the data should be regarded as a conservative estimate of the entanglement that can be effectively harnessed under the present level of system stability.
@@
We also note that the rectangular spectral filter used here is not optimal. Tailored mode shaping could increase the opto-mechanical fidelity and enhance the observable entanglement [7].
@@
The results discussed so far were obtained for detunings satisfying $\Delta_{A}\simeq\Delta_{B}\simeq\Omega_{\mathrm{b}}$. This condition, however, is not essential, and a finite degree of entanglement is present even when it is relaxed. Indeed, we repeated the experiment while varying the detuning, keeping only the condition $\Delta_{A}\simeq\Delta_{B}$. Fig. 4 shows the symplectic eigenvalue $\nu_{-}$ derived from the covariance matrix $\mathbf{V}$ reconstructed directly from the measured correlation matrix $\mathbf{A}$, together with the corresponding values obtained by computing $\mathbf{V}$ with the optomechanical model using parameters fitted to the same experimental data.
Entanglement persists over a detuning range exceeding 40 kHz, consistent with the expectation that it is maintained within a bandwidth on the order of the cavity linewidth $\kappa$. This robustness implies that stable opto-mechanical entanglement can be achieved without fine control of the detuning.
@@
Figure 4: Symplectic eigenvalue versus cavity detuning. Experimental data points shown by blue markers with error bars (SEM over about 10 independent data acquisitions) are obtained from the covariance matrix directly reconstructed from the measured correlations. Green markers are extracted instead from the same experimental data through the optomechanical model, using the parameters that best fit the heterodyne spectra. In this case the uncertainty is dominated by the systematic error on the detection efficiency $\eta$. In these measurements, we set $\Delta_{A}\simeq\Delta_{B}=|\Delta|$. Intracavity entanglement is inferred from the symplectic eigenvalues associated to the covariance matrix calculated, within the same model and parameter set, for the joint state of the motion along an optimized direction and the intracavity $B$ field (red points, with error bars again set by the systematic uncertainty on $\eta$). The green and red shaded regions are obtained by taking the extremal model predictions generated by the full set of parameters extracted from all data sets (at all detunings). These bands therefore illustrate the overall detuning dependence of the entanglement in our system.
@@
An important and nontrivial question is whether entanglement is also present inside the cavity, between the intracavity mode of field $B$ and the mechanical oscillator. This is reasonable but not obvious, since the intracavity field mode is uniquely determined by the cavity boundary conditions and cannot be tailored as in the case of the propagating field. Although intracavity observables cannot be accessed directly, the same optomechanical model that quantitatively reproduces the measured output correlations allows us to reconstruct the corresponding intracavity covariance matrix from the fitted parameters. A crucial issue is the motion along the tweezer axis (the $z$ mode), whose mechanical frequency is $\Omega_{z}/2\pi=19.3\,$kHz. This motion is coupled to the cavity field due to a small angle between the cavity axis and the tweezer transverse plane. As a consequence, the $z$ mode produces a peak also in the spectrum of the bright mode through the cavity-mediated coupling, slightly increasing its variance. We have indeed added this contribution in the entries $\mathrm{V}_{33}$ and $\mathrm{V}_{44}$ before computing the values of $\nu_{-}$ here reported [18]. The $z$ motion has no effect on the propagating optical mode, because the filter $\xi$ removes the spectral region around the $z$ mode resonance. In contrast, the $z$ mode peak is not negligible in the spectra of the cavity optical modes and, in fact, it destroys the entanglement with the transverse mechanical bright mode considered so far. Nevertheless, optomechanical entanglement is partially recovered by considering a slightly tilted direction of motion, thereby including a projection of the tweezer axis. This enhances the optomechanical correlations in the spectral region of the $z$ mode. The symplectic eigenvalue $\nu_{-}$ associated with the covariance matrix describing the cavity $B$ mode and the optimized motion direction, shown in Fig. 4 as a function of the detuning, confirms the presence of intracavity entanglement, albeit at a reduced level compared to that of the propagating mode. The minimum inferred value of $\nu_{-}$ for the intracavity variables is $\nu_{-}=0.976\pm 0.007$, corresponding to a logarithmic negativity of $E_{N}=0.035\pm 0.010$, and occurs at a detuning of $|\Delta|/2\pi=137$ kHz.
@@
In conclusion, we demonstrate quantum entanglement between the center-of-mass motion of a levitated nanosphere and a propagating optical field. By combining coherent scattering in an optical cavity with high efficiency heterodyne detection, we reconstruct the full set of optical–mechanical correlations and observe the violation of separability bounds under experimentally accessible conditions. These results establish levitated optomechanical systems as a versatile platform for generating and distributing light–matter entanglement at room temperature, without ultra-cryogenic environments or stringent parameter tuning, providing a significant contribution to macroscopic quantum technologies. Beyond center-of-mass motion, additional mechanical degrees of freedom — including rotational and librational modes [39, 24, 12, 40] — provide further opportunities for multimode quantum control. Combined with the prospect of trapping and simultaneously cooling multiple levitated nanospheres in a single optical cavity [50, 38], this architecture could enable multipartite entanglement mediated by a shared optical field [9]. More broadly, the ability to prepare nonclassical states of increasingly massive objects may open the way to experimental tests of the possible quantum nature of gravity [6, 30] and conjectured modifications of quantum mechanics at macroscopic scales.
@@ Acknowledgments
The authors would like to acknowledge useful discussions with D. Vitali, P. F. Barker and U. Delić. We acknowledge financial support from PNRR MUR Project No. PE0000023-NQSTI and by the European Commission-EU under the Infrastructure I-PHOQS “Integrated Infrastructure Initiative in Photonic and Quantum Sciences ” [IR0000016, ID D2B8D520, CUP D2B8D520].